OpenStudy (anonymous):

The down payment for the lot you buy is $25,000. If you have only $9500, what percent of the down payment do you still need?

OpenStudy (anonymous):

set it up as a proportion, PART over WHOLE equals the money you own over the the money you owe\[\frac{ PART }{ WHOLE }\] \[= \frac{ $OWN }{ $OWE}\] s so that would mean \[\frac{ 9500 }{ 25000 }= \frac{ x }{ 100 }\]